# Interpolating many-body wave functions for accelerated molecular dynamics on near-exact electronic surfaces

### v1

@@ Introduction
It was found important to develop an active learning scheme for the selection of appropriate atomic configurations to include in the training data for rapid convergence. In Ref. [mejutozaera2023quantum], the energy variance was motivated as an appropriate measure for the inclusion of data points, however this is impractical in the current lower-scaling scheme as it would require the evaluation of higher-body transition density matrices between training states. Instead, we consider the addition of training points which will maximize the improvement in the MD trajectories while respecting the invariances in the model predictions. This is performed by selecting the point on the trajectory where the Hamiltonian operator in the SAO basis $K_{ijkl}(\mathbf{R})$ has changed most (in a least squares sense) compared to the Hamiltonians used in the current training data set. Since the probability amplitudes are uniquely defined by this Hamiltonian, it is a suitable measure for the addition of new data points. Furthermore, due to the variationality of the method, it is guaranteed that the potential energy with the enlarged training data will be equal or lower to the previous predictions, across the whole trajectory. This can therefore be used as a rigorous metric for the systematic convergence of the potential energy surface for the MD, with more details in the Methods section. We consider the potential energy surface over the whole MD simulation fully converged when the maximum reduction in energy for any point over the whole trajectory is less than $1\,mE_{\mathrm{h}}$ for two consecutive increases in the data set size.
@@
Figure 4: Molecular dynamics of a $30$ atom hydrogen chain in a STO-6G basis from an initially symmetrically stretched geometry. The panels report the distance between two of the hydrogen atoms (top panel: first and last atom in the chain, bottom panel: first and second atom) against the simulation time, showing that the first two hydrogen atoms form a stable vibrating dimer while the chain lengthens. The trajectories were obtained from the eigenvector continuation with $N=47$ (blue, solid), $N=10$ (green, dashed) and $N=2$ (red, dotted) DMRG data points, together with the trajectories from DMRG (black, dashed), Hartree-Fock (grey, dashed) and density functional theory with a PBE exchange correlation (grey, dash-dotted) potential energy surfaces. Additional snapshots depict the initial and final hydrogen chain arrangements obtained from the converged eigenvector continuation ($N=47$).
@@
A semi-infinite symmetric one-dimensional chain of hydrogen atoms has emerged as a paradigmatic benchmark system of strongly correlated electronic structure in recent years, as a platform towards larger ab initio and extended systems. Almost all modern electronic structure methods have been applied to the system with varying success, and it has motivated further developments in both theory and understanding of its unexpectedly rich phase diagram [PhysRevX.7.031059, PhysRevX.10.031058]. While the symmetric stretch of this system has been considered extensively via single-point electronic structure, its full dynamics at this level has not. In Fig. 4, we release the atoms to dynamically move on a tightly-converged ground state surface of the DMRG-trained continuation scheme, starting from a $\sim$10symmetric stretching of thirty atoms equally from the symmetric equilibrium structure. We find that along with the vibrations of the bonds, the atoms rapidly dimerize and separate, with the overall length of the chain increasing approximately linearly with time. We are able to converge the dynamics of this dimerization and dissociation exactly (albeit in a minimal basis) with only a small number of DMRG calculations. We note that in comparison, DFT-based AI-BOMD significantly underestimates the rate of dimerization of the chain, while Hartree–Fock theory conversely results in a bond for the hydrogen dimers which is too stiff, demonstrating the importance of an accurate treatment of the electronic correlations in the dynamics.
Realistic chemistry on the exact electronic surface
@@
We consider the feasibility of converging faithful reactive chemistry on a near-exact potential energy surface for the gas-phase dynamics of a Zundel cation, comprising a water molecule and hydronium ion - a system whose intricate potential energy surface poses a challenging test case for novel numerical techniques, yet is particularly important in the understanding of proton diffusion in aqueous solution [Huang2005AbIP, McCoy2005FulldimensionalVC, KondatiNatarajan2015RepresentingTP, Heindel2018BenchmarkES, 10.1063/1.4941091]. We start the simulation from a high-energy configuration away from the traditional Grotthuss mechanism, where the additional proton bridges the two ions, interrupting the hydrogen bond as shown in Fig. 5. We rigorously converge the potential energy surface of the MD with the continuation scheme based on DMRG training data, which required 84 training points to cover the $120\,fs$ of the simulation (see extended data for evidence of this convergence with training data over the trajectory). Due to the fact that an explicit representation of the electronic state is retained over the trajectory, we also extract non-energetic electronic properties of the system over time. We use this to consider the evolution of the Mulliken charge as the electron density is redistributed around the system in response to the atomic motion, beyond the physics considered in traditional polarizable force fields.
@@
Figure 5: Predicted Mulliken charges for the hydronium (top panel) and water (bottom panel) sub-units from an MD simulation of the reaction from an eigenvector continuation with $N=40$ (green, dashed) and $N=84$ (blue, solid) DMRG data points. The system uses a 6-31G basis, with snapshots depicting the evolution of the molecular geometry at four evenly spaced times. Reference charges from simulations with DFT with a B3LYP exchange correlation (dash-dotted) and HF (dotted) are included for comparison.
@@
The positive charge is initially fairly evenly distributed amongst the sub-units, with the anticipation being that the water would rotate to adopt a lower-energy configuration. However, on the DMRG interpolated potential energy surface we find that before this is able to occur, a (neutral) hydrogen is ejected from the system leaving a bound hydronium and hydroxide radical system. The charges on the sub-units of this reaction show the redistribution of charge as the hydrogen oscillates a number of times before its eventual ejection from the system. This behavior is not seen on the more approximate HF or DFT electronic surfaces where the additional hydrogen remains bound over timescales enabling the water to rotate its orientation, with the HF charge distribution substantially in error even in the initial state.
@@
While these results describe the converged dynamics of a single trajectory in a relatively restricted 6-31G basis, more consideration of a statistical ensemble of initial states, as well as potentially basis size, thermalization at finite temperature, nuclear quantum and solvent effects may be needed before predictions as to the nature of this physical reaction can be given with confidence [10.1063/1.4941091]. However, the fact that qualitative changes in dynamical behavior already result from the quality of the treatment of electronic correlation effects in the determination of the potential energy surface underlines the importance of a robust and systematically improvable approach to compute this surface. The eigenvector continuation acceleration allows computation of this surface with high-level quantum chemical methods, and extends their scope to enable them to access timescales of atomic dynamics with provable convergence.
Outlook
@@
We develop a practical approach for eigenvector continuation of many-body electronic wave functions in ab initio settings. In contrast to the traditional paradigm of machine-learning force fields from training energies, this considers the interpolation of accurate wave functions across chemical space, from which all electronic properties as well as atomic forces can be efficiently computed via a variational ansatz, avoiding the exponential complexity of the many-body states themselves. Using the scheme to converge the potential surface for molecular dynamics, we find examples of qualitatively different behavior to state-of-the-art techniques, demonstrating the importance of systematically converging the electronic structure across the timescales.
@@ Methods
Interpolating across ab initio potential energy surfaces
@@
with Fermionic creation and annihilation operators, $\hat{c}^{\dagger}$ and $\hat{c}$ acting on the orbitals. The one-electron terms, $h^{(1)}_{ij}(\mathbf{R})$, are matrix elements of the electron-nuclear and electronic kinetic operators, while the electron-electron repulsion integrals are
@@
$=\langle ij|kl\rangle.$ (8)
A convenient reduced two-body Hamiltonian which subsumes the one-body into the two-body term can be written as
$K_{ijkl}=\frac{1}{2}\langle ij|kl\rangle+\frac{1}{2(N_{\mathrm{elec}}-1)}(\delta_{jl}h^{(1)}_{ik}+\delta_{ik}h^{(1)}_{jl}),$ (9)
with $E_{\mathrm{nuc}}(\mathbf{R})$ the classical nuclear-nuclear repulsion term [PhysRevA.57.4219].
@@
The continuation then proceeds according to the scheme outlined in the main text, with the evaluation of the transition two-body density matrices (t-2RDMs) and overlaps between the training points in their SAO representations. Of particular importance for molecular dynamics is the evaluation of analytic forces at each test geometry, which can be achieved via application of the Hellmann-Feynman theorem, which is simple due to the fully optimized variational nature of the interpolated states in the geometry-independent basis [PhysRev.56.340, 10.1063/1.4927594]. This therefore only required the derivatives of the electron integrals in the AO basis [https://doi.org/10.1002/jcc.23981, https://doi.org/10.1002/wcms.1340, 10.1063/5.0006074], as well as the derivative of the transformation from the atomic orbitals to the SAOs with respect to nuclear positions, which we evaluate via first order perturbation theory [bamieh2022tutorial]. The specifics of this evaluation is given in the Supplementary information.
@@
Firstly, we consider the compression of the training wave functions in the form of Matrix Product States (MPS), optimized via the density matrix renormalization group (DMRG) algorithm [PhysRevLett.69.2863]. For this, we used the spin-adapted implementation from the block2 library [10.1063/1.3695642, 10.1063/5.0050902], working directly in the Fock space defined by the SAOs. We optimize the training states with a schedule for exponentially increasing bond dimension (a factor of $1.8$ per increase) and decreasing noise in the MPS, a standard practice for stable ab initio DMRG [10.1063/1.4905329], terminating when the difference of the energy upon fully relaxing the state at an increased bond dimension is less than a specified threshold, $\epsilon$. For all presented results, we set $\epsilon=10^{-3}\,E_{\mathrm{h}}$ and start the MPS with an initial bond dimension of $34$, giving training data confidently below the accepted ‘chemical accuracy’. For the reference data for the hydrogen chain evolution (Fig. (4)), we set the tolerance to $\epsilon=10^{-5}\,E_{h}$, and starting MPS bond dimension to $61$.
@@
While this state can be straightforwardly optimized within a ‘CAS-CI’ scheme implemented in the PySCF package [https://doi.org/10.1002/wcms.1340, 10.1063/5.0006074], we also require the evaluation of the overlap and the t-2RDMs between training states in their SAO basis, while the state is defined (and optimized) in a mean-field canonical basis. Therefore, it is necessary to rotate these many-body states into their respective SAO bases before the t-2RDMs and overlaps computed. We show this for the t-2RDM as
@@
where $C^{(a/b)}_{\mathbf{n}/\mathbf{n}^{\prime}}$ are the CAS-CI probability amplitudes of the active spaces. This single-particle unitary transformation $\hat{U}^{\dagger}_{\mathbf{R}}$ can be formed as
@@
The inner products $\langle\mathbf{n}|\hat{U}^{\dagger}_{\mathbf{R}^{(a)}}\hat{c}^{\dagger}_{i}\hat{c}^{\dagger}_{j}\hat{c}_{k}\hat{c}_{l}\hat{U}_{\mathbf{R}^{(b)}}|\mathbf{n}^{\prime}\rangle$ from Eq. (14) can be identified as a matrix element between two different non-orthogonal Slater determinants [doi:10.1021/acs.jctc.5b01148]. The efficient evaluation of such overlaps between different non-orthogonal Slater determinants is discussed in Refs. [10.1063/5.0045442, 10.1063/5.0122094]. We utilize the libgnme package, together with its python interface pygnme, to evaluate the overlaps and t-2RDMs between CAS states for the continuation in the SAO basis. Due to the non-orthogonality of the different CAS spaces, the double contraction of Eq. (14) results in a cost scaling quadratically in the size of the active space, thus more expensive than the evaluation of expectation values of a single point CAS state, however this cost could be reduced in the future by rotating to an intermediate basis representing the co-domain of the occupied CAS orbitals in a pair of CAS-CI training states.
We note here that the eigenvector continuation scheme based on approximate solvers (especially CAS-CI) can result in inferred energies at training points which are variationally lower than those of the corresponding training energy at the same geometry. This is because it can mix contributions to the test state from other training states, with the resulting linear combination of approximate states lower in energy, resulting in an improved estimate for the energy at a test geometry than its corresponding training energy. Fig. 6 of the extended data section exemplifies this behaviour, indicating lower predicted energies from the continuation of CAS states as compared to the reference CAS energies along the trajectory of a vibrating water molecule.
@@
The Born–Oppenheimer molecular dynamics trajectories were computed in vacuum based on a microcanonical (NVE) ensemble using the Velocity-Verlet integration implemented in PySCF [https://doi.org/10.1002/wcms.1340, 10.1063/5.0006074, 10.1063/1.442716, marx2000ab]. This solves Newton’s equation for the position of each nucleus,
$m_{i}\frac{\mathrm{d}^{2}\mathbf{R}_{i}}{\mathrm{d}t^{2}}=\mathbf{F}_{i}(\mathbf{R}),$ (16)
where $\mathbf{R}_{i}$ is the position of the nucleus indexed by $i$, with mass $m_{i}$ and associated force component $\mathbf{F}_{i}$ according to the analytic nuclear gradients derived for the eigenvector continuation in the supplementary information. In all tests, we initialized the nuclei at rest, and chose a fixed timestep of $\delta t=5\,a.u\approx 0.121\,fs$ for the integration.
@@
To gauge the systematic convergence of the MD trajectory, we track the variational lowering (and hence improvement) of the potential energy surface as the dataset is enlarged. This is done by comparing the smaller and larger PES along the same trajectory corresponding to the larger data set. Due to the variationality of the method, it is guaranteed that the potential energy inference with the larger dataset will be lower or equal to the predictions with the smaller dataset, and we use the difference between the predicted energies as a convergence measure. In our applications, we terminate the simulation when the predicted energy with the enlarged data set stays within a tolerance of $\epsilon=10^{-3}\,E_{\mathrm{h}}$ along the full MD trajectory for two iterations in a row. Examples of this convergence are shown in the extended data.
@@
The code and inputs to fully reproduce the numerical experiments of this work can be found at https://github.com/BoothGroup/evcont. Furthermore, we provide the videos of the molecular dynamics trajectories presented in this work at https://doi.org/10.5281/zenodo.10658975.

### v2

@@ Abstract
It was found important to develop an active learning scheme for the selection of appropriate atomic configurations to include in the training data for rapid convergence. In Ref. [mejutozaera2023quantum], the energy variance was motivated as an appropriate measure for the inclusion of data points, however this is impractical in the current lower-scaling scheme as it would require the evaluation of higher-body transition density matrices between training states. Instead, we consider the addition of training points which will maximize the improvement in the MD trajectories while respecting the invariances in the model predictions. This is performed by selecting the point on the trajectory where the Hamiltonian operator in the SAO basis, $K_{ijkl}(\mathbf{R})$, has changed most (in a least squares sense) compared to the Hamiltonians employed to generate the current training data set. Since the probability amplitudes are uniquely defined by this Hamiltonian, it is a suitable measure for the addition of new data points. Furthermore, due to the variationality of the method, it is guaranteed that the potential energy with the enlarged training data will be equal or lower to the previous predictions, across the whole trajectory. This can therefore be used as a rigorous metric for the systematic convergence of the potential energy surface for the MD, with more details in the Methods section. We consider the potential energy surface over the whole MD simulation fully converged when the maximum reduction in energy for any point over the whole trajectory is less than $1\,mE_{\mathrm{h}}$ for two consecutive increases in the data set size.
@@
Figure 4: Molecular dynamics of a $30$ atom hydrogen chain in a STO-6G basis from an initially symmetrically stretched geometry. The panels report the distance between two of the hydrogen atoms (top panel: first and last atom in the chain, bottom panel: first and second atom) against the simulation time, showing that the first two hydrogen atoms form a stable vibrating dimer while the chain lengthens. The trajectories were obtained from the eigenvector continuation with $N=47$ (blue), $N=10$ (green) and $N=2$ (red) DMRG data points, together with the trajectories from DMRG (black, dashed), Hartree-Fock (brown, dotted) and density functional theory with a PBE exchange correlation (light blue, dash-dotted) potential energy surfaces. Additional snapshots depict the initial and final hydrogen chain arrangements obtained from the converged eigenvector continuation ($N=47$).
@@
A semi-infinite symmetric one-dimensional chain of hydrogen atoms has emerged as a paradigmatic benchmark system of strongly correlated electronic structure in recent years, as a platform towards larger ab initio and extended systems. Almost all modern electronic structure methods have been applied to the system with varying success, and it has motivated further developments in both theory and understanding of its unexpectedly rich phase diagram [PhysRevX.7.031059, PhysRevX.10.031058]. While the symmetric stretch of this system has been considered extensively via single-point electronic structure, its full dynamics at this level has not. In Fig. 4, we release the atoms to dynamically move on a tightly-converged ground state surface of the DMRG-trained continuation scheme, starting from a $\sim$10symmetric stretching of thirty atoms equally from the symmetric equilibrium structure. We find that along with the vibrations of the bonds, the atoms rapidly dimerize and separate, with the overall length of the chain increasing approximately linearly with time. We are able to converge the dynamics of this dimerization and dissociation (albeit in a minimal basis) to the equivalent explicit DMRG AI-BOMD with only a small number of single-point training DMRG calculations. We note that in comparison, DFT-based AI-BOMD significantly underestimates the rate of dimerization of the chain, while Hartree–Fock theory conversely results in a bond for the hydrogen dimers which is too stiff, demonstrating the importance of an accurate treatment of the electronic correlations in the dynamics.
Towards chemical accuracy for realistic (thermo)chemistry
We consider the feasibility of converging faithful thermodynamic quantities and reactive chemistry on a near-exact potential energy surface for the gas-phase dynamics of a Zundel cation, comprising a water molecule and hydronium ion - a system whose intricate potential energy surface poses a challenging test case for novel numerical techniques, yet is particularly important for the understanding of proton diffusion in aqueous solution [D2SC03189B, Huang2005AbIP, McCoy2005FulldimensionalVC, KondatiNatarajan2015RepresentingTP, Heindel2018BenchmarkES, 10.1063/1.4941091]. We first consider a statistical ensemble of $500$ different trajectories, starting from the same geometry (taken from Ref. [10.1063/1.1834500]), and sampling initial velocities from a Maxwell–Boltzmann distribution at $298.15\,\mathrm{K}$. The BOMD was propagated under NVT conditions to thermalize according to a Berendsen integration scheme [berendsen1984molecular]. We consider $N=60,80$ and $100$ single-point DMRG training configurations to observe the convergence of the thermodynamically equilibrated properties in a 6-31G basis. Each ensemble of trajectories at one of these training number involved $5\times 10^{6}$ potential energy and force evaluations, which would be out of reach with a brute-force DMRG approach, but required a relatively modest 7,500 CPU hours for the propagation of the full ensemble. Nevertheless, we can explicitly verify convergence to the accuracy of the underlying DMRG by a validation of the ‘test error’ via additional DMRG calculations for sampled geometries along the trajectories. The achieved test error, shown in the extended data, demonstrates that the PES is well below chemical accuracy of the exact potential energy surface within the employed basis as the thermal equilibrium is approached, reaching relative correlation energy errors below that of both CCSD and CCSD(T) – the ‘gold standard’ of quantum chemistry [RevModPhys.79.291].
Figure 5: Thermalization to $298.15\mathrm{K}$ of the trajectory-averaged distance between the central hydrogen and the oxygen atom of the Zundel cation. The main plot shows the mean distance as a function of propagation time obtained by interpolating from $N=60,80,100$ DMRG calculations, as well as DFT (with CAM-B3LYP [YANAI200451] and PBE exchange-correlation functionals) and CCSD trajectories. The mean corresponds to a running average of the distance between the atoms with a window of $100$ timesteps ($\approx 60.4\,\mathrm{fs}$), and averaging over $500$ independent trajectories and both oxygen atoms. Each emsemble of trajectories required $5$ million energy and force calculations. Inset shows the thermalized radial distribution function of the oxygen from the central hydrogen, using a Gaussian smearing of individual data points in a kernel density analysis [chen2017tutorial, scikit-learn], with a bandwidth of $\sigma=0.0025\,$Å.
Figure 5 shows this thermalization in the average distance between the central hydrogen atom and the two oxygen atoms in the explored Zundel configurations. We find this statistically equilibrated distance to be converging to a slightly shorter length than CCSD as the number of training configurations is increased. An accurate description of this multi-center bond is key for the Grotthuss mechanism of proton transfer. The differences in these quantities are in stark contrast to the much shorter distances predicted by DFT MD simulations with two widely used exchange–correlation functionals, which indicate a more localized central hydrogen. We can observe this in the radial distribution function of the equilibrated configurations (inset) where the distribution is far flatter than the DFT methods, indicating an increased delocalization of the hydrogen between the water subunits. This is further corroborated by considering the magnitude of the dipole moment from the thermalized ensemble (see extended data), which we find decreases as the level of theory is increased from DFT to CCSD to the DMRG-interpolated configurations, indicating a preference for more symmetric distributions where the central hydrogen is delocalized and less bound at any instant to an individual oxygen atom.
The verifiably high-accuracy interpolation coupled with the high-accuracy DMRG training allows for validation in the use of CCSD for this system, with thermalized expectation values qualitatively in agreement. This is due to the lack of strong correlation in the explored molecular configurations. However, a significant advantage of this framework is the ability to also reliably converge the PES over the full phase space, including strongly correlated atomic configurations further from equilibrium where CCSD is unreliable and will potentially fail, including bond-breaking and transition state geometries.
@@
To consider this scenario, we also propagate a single high-energy trajectory within an NVE ensemble far from the Grotthuss mechanism dynamics, where the additional proton is inserted between the ions, interrupting the traditional hydrogen bond framework with a four-atom bridging bond as shown in the initial snapshot of Fig. 6. Increasing the number of DMRG training points to $N=84$, we are able to observe convergence in the specific short-time MD trajectory over the $120\,\mathrm{fs}$ of the simulation (see extended data for evidence of this convergence with training data over the trajectory). Due to the fact that an explicit representation of the electronic state is retained over the trajectory, we also extract non-energetic electronic properties of the system over time. We use this to consider the evolution of the Mulliken charge as the electron density is redistributed around the system in response to the atomic motion, beyond the physics considered in traditional polarizable force fields.
@@
Figure 6: Predicted Mulliken charges for the hydronium (top panel) and water (bottom panel) sub-units from an MD simulation of the reaction from an eigenvector continuation with $N=40$ (green) and $N=84$ (blue) DMRG data points. The system uses a 6-31G basis, with snapshots depicting the evolution of the molecular geometry at four evenly spaced times. Reference charges from simulations with DFT with a B3LYP exchange correlation (ligh blue, dash-dotted) and HF (brown, dotted) are included for comparison.
@@
The positive charge is initially fairly evenly distributed amongst the water monomers, with the anticipation being that the water would rotate to adopt a lower-energy configuration. However, on the DMRG interpolated PES we find that before this is able to occur, a (neutral) hydrogen is ejected from the system leaving a bound hydronium and hydroxide radical system. The charges on the sub-units of this reaction show the redistribution of charge as the hydrogen oscillates a number of times before its eventual ejection from the system. This behavior is not seen on the more approximate HF or DFT electronic surfaces where the additional hydrogen remains bound over timescales enabling the water to rotate its orientation, with the HF charge distribution substantially in error even in the initial state. Comparing to CCSD across the $N=84$ trajectory, we find $\approx 40$ atomic configurations visited result in the CCSD energy diverging due to the presence of strong correlation effects, underlining the unreliability of the method for MD in these more unusual atomic conformations and transition states where strongly correlated electronic structure is found. This behavior is discussed further in the extended data, but underlines the applicability of the continuation across the phase space of the MD and the potential to describe dissociative dynamics [D4FD00062E].
@@
While these results demonstrate an effective acceleration scheme to converge the energy surface of this system to that of high accuracy methods, more consideration of the effects of basis size, nuclear quantum and solvent effects may be needed before predictions as to the nature of this physical reaction can be given with confidence [10.1063/1.4941091, C4CP05192K]. However, the fact that qualitative changes in dynamical behavior already result from the quality of the treatment of electronic correlation effects in the determination of the potential energy surface underlines the importance of a robust and systematically improvable approach to this electronic structure. The eigenvector continuation acceleration allows computation of this surface with high-level quantum chemical methods, and extends their scope to enable them to access timescales of atomic dynamics with provable convergence.
Conclusion
@@
We develop a practical approach for eigenvector continuation of many-body electronic wave functions in ab initio settings. In contrast to the traditional paradigm of machine-learning force fields from training energies, this considers the interpolation of accurate wave functions across the space of structural changes, from which all electronic properties as well as atomic forces can be efficiently computed via a variational ansatz, avoiding the exponential complexity of the many-body states themselves. Using the scheme to converge the potential surface for molecular dynamics, we find examples of qualitatively different behavior to state-of-the-art techniques, demonstrating the importance of systematically converging the electronic structure across the timescales.
@@ Methods
Interpolating across the ab initio potential energy surface
@@
with Fermionic creation and annihilation operators, $\hat{c}^{\dagger}$ and $\hat{c}$ acting on the orbitals, and $E_{\mathrm{nuc}}(\mathbf{R})$ the classical nuclear-nuclear repulsion energy. The one-electron terms, $h^{(1)}_{ij}(\mathbf{R})$, are matrix elements of the electron-nuclear and electronic kinetic operators, while the electron-electron repulsion integrals are
@@
$=\langle ij|kl\rangle{(\mathbf{R})}.$ (8)
A convenient reduced two-body Hamiltonian which subsumes the one-body into the two-body term can be written as [PhysRevA.57.4219]
$K_{ijkl}{(\mathbf{R})}$ $=\frac{1}{2}\langle ij|kl\rangle{(\mathbf{R})}$ (9)
$+\frac{1}{2(N_{\mathrm{elec}}-1)}(\delta_{jl}h^{(1)}_{ik}{(\mathbf{R})}+\delta_{ik}h^{(1)}_{jl}{(\mathbf{R})}).$
@@
The continuation then proceeds according to the scheme outlined in the main text, with the evaluation of the transition two-body density matrices (t-2RDMs) and overlaps between the training points in their SAO representations. Of particular importance for molecular dynamics is the evaluation of analytic forces at each test geometry, which is simplified due to the lack of response contributions from the many-body basis and the fully optimized variational nature of the interpolated states in the geometry-independent basis [PhysRev.56.340, 10.1063/1.4927594]. This therefore only required the derivatives of the electron integrals in the AO basis [https://doi.org/10.1002/jcc.23981, https://doi.org/10.1002/wcms.1340, 10.1063/5.0006074], as well as the derivative of the transformation from the atomic orbitals to the SAOs with respect to nuclear positions (a ‘Pulay force’ [doi:10.1080/00268976900100941]), which we evaluate via first order perturbation theory [bamieh2022tutorial]. The specifics of this evaluation is given in the Supplementary information.
@@
Firstly, we consider the compression of the training wave functions in the form of Matrix Product States (MPS), optimized via the density matrix renormalization group (DMRG) algorithm [PhysRevLett.69.2863]. For this, we used the spin-adapted implementation from the block2 library [Block2, 10.1063/5.0050902], working directly in the Fock space defined by the SAOs. We optimize the training states with a schedule for exponentially increasing bond dimension (a factor of $1.8$ per increase) and decreasing noise in the MPS, a standard practice for stable ab initio DMRG [10.1063/1.4905329], terminating when the difference of the energy upon fully relaxing the state at an increased bond dimension is less than a specified threshold, $\epsilon$. For all presented results, we set $\epsilon=10^{-3}\,E_{\mathrm{h}}$ and start the MPS with an initial bond dimension of $34$, giving training data confidently below the accepted ‘chemical accuracy’. For the reference data for the hydrogen chain evolution (Fig. (4)), we set the tolerance to $\epsilon=10^{-5}\,E_{\mathrm{h}}$, and starting MPS bond dimension to $61$.
@@
While this state can be straightforwardly optimized within a ‘CASCI’ scheme implemented in the PySCF package [https://doi.org/10.1002/wcms.1340, 10.1063/5.0006074], we also require the evaluation of the overlap and the t-2RDMs between training states in their SAO basis, while the state is defined (and optimized) in a mean-field canonical basis. Therefore, it is necessary to rotate these many-body states into their respective SAO bases before the t-2RDMs and overlaps computed. We show this for the t-2RDM as
@@
where $C^{(a/b)}_{\mathbf{n}/\mathbf{n}^{\prime}}$ are the CASCI probability amplitudes of the active spaces. This single-particle unitary transformation $\hat{U}^{\dagger}_{\mathbf{R}}$ can be formed as
@@
The inner products $\langle\mathbf{n}|\hat{U}^{\dagger}_{\mathbf{R}^{(a)}}\hat{c}^{\dagger}_{i}\hat{c}^{\dagger}_{j}\hat{c}_{k}\hat{c}_{l}\hat{U}_{\mathbf{R}^{(b)}}|\mathbf{n}^{\prime}\rangle$ from Eq. (14) can be identified as a matrix element between two different non-orthogonal Slater determinants [doi:10.1021/acs.jctc.5b01148]. The efficient evaluation of such overlaps between different non-orthogonal Slater determinants is discussed in Refs. [10.1063/5.0045442, 10.1063/5.0122094]. We utilize the libgnme package, together with its python interface pygnme, to evaluate the overlaps and t-2RDMs between CAS states for the continuation in the SAO basis. Due to the non-orthogonality of the different CAS spaces, the double contraction of Eq. (14) results in a cost scaling quadratically in the size of the active space, thus more expensive than the evaluation of expectation values of a single point CAS state, however this cost could be reduced in the future by rotating to an intermediate basis representing the co-domain of the occupied CAS orbitals in a pair of CASCI training states.
We note here that the eigenvector continuation scheme based on approximate solvers (especially CASCI) can result in inferred energies at training points which are variationally lower than those of the corresponding training energy at the same geometry. This is because it can mix contributions to the test state from other training states, with the resulting linear combination of approximate states lower in energy, resulting in an improved estimate for the energy at a test geometry than its corresponding training energy. Fig. 7 of the extended data section exemplifies this behaviour, indicating lower predicted energies from the continuation of CAS states as compared to the reference CAS energies along the trajectory of a vibrating water molecule. In addition, discontinuities in CASCI potentials due to state crossings in the active space are possible and plague active space methods in general for complex trajectories. Due to the linearity of the ansatz in the interpolated energy surface, these discontinuities are necessarily removed, with an example for the water vibration of this phenomena shown in the extended data.
@@
The single-trajectory Born–Oppenheimer molecular dynamics of Figs. 3, 4 and 6 were computed in vacuum based on a microcanonical (NVE) ensemble using the Velocity-Verlet integration implemented in PySCF [https://doi.org/10.1002/wcms.1340, 10.1063/5.0006074, 10.1063/1.442716, marx2000ab], according to the analytic nuclear gradients derived for the eigenvector continuation in the supplementary information. The nuclei in these simulations were initialized at rest, and chose a fixed timestep of $\delta t=5\,\mathrm{a.u.}\approx 0.121\,\mathrm{fs}$ for the integration.
[…]
@@
To gauge the systematic convergence of the NVE MD single-shot trajectories, we can track the variational lowering (and hence improvement) of the potential energy surface as the dataset is enlarged. This is done by comparing the PES from the two data set sizes along the same trajectory corresponding to the larger of the two data sets. Exploiting the variationality of the method, it is guaranteed that the potential energy inference with the larger dataset will be lower or equal to the predictions with the smaller dataset, and we use the difference between the predicted energies as a convergence measure. In our applications, we terminate the simulation when the predicted energy with the enlarged data set stays within a tolerance of $\epsilon=10^{-3}\,E_{\mathrm{h}}$ along the full MD trajectory for two iterations in a row. Examples of this convergence are shown in the extended data.
To manage the increased data volume when generating the statistical canonical ensemble of trajectories for the NVT Zundel cation results of Fig. 5, we use a somewhat coarser scheme to select the training configurations. We start with just the initial configuration in the training set, and randomly sub-sample $100$ trajectories from the ensemble of $500$ trajectories generated by the prior CAM-B3LYP DFT dynamics. For all timesteps of these $100$ trajectories, the Hamiltonian distance metric of Eq. 16 is computed and the $19$ geometries with the largest value of this metric are identified for inclusion in an enlarged training data set. The continuation scheme is then run for $500$ NVT trajectories with these $20$ training points. We compute the Hamiltonian metric along the full path of a new random selection of $100$ of these inferred trajectories in order to identify a further set of $20$ geometries to perform explicit DMRG calculations to iteratively enlarge the training data set until the desired size is reached. This training set is taken to be the same for all trajectories in an ensemble. It should be stressed that only the first batch of $19$ geometries are taken from the DFT-derived trajectories, after which subsequent batches of training geometries are found self-consistently to ensure a systematically reducing bias due from the DFT paths.
@@
The code and inputs to fully reproduce the numerical experiments of this work can be found at https://github.com/BoothGroup/evcont. Furthermore, we provide the videos of the isolated NVE molecular dynamics trajectories presented in this work at https://doi.org/10.5281/zenodo.10658975.